# Improving the (approximate) sequential probability ratio test by avoiding overshoot

### v1

@@ Abstract
The sequential probability ratio test (SPRT) by Wald (1945) is a cornerstone of sequential analysis. Based on desired type-I, II error levels $\alpha,\beta\in(0,1)$, it stops when the likelihood ratio statistic crosses certain upper and lower thresholds, guaranteeing optimality of the expected sample size. However, these thresholds are not closed form and the test is often applied with approximate thresholds $(1-\beta)/\alpha$ and $\beta/(1-\alpha)$ (approximate SPRT). When $\beta>0$, this neither guarantees type I,II error control at $\alpha,\beta$ nor optimality. When $\beta=0$ (power-one SPRT), it guarantees type I error control at $\alpha$ that is in general conservative, and thus not optimal. The looseness in both cases is caused by overshoot: the test statistic overshoots the thresholds at the stopping time. One standard way to address this is to calculate the right thresholds numerically, but many papers and software packages do not do this. In this paper, we describe a different way to improve the approximate SPRT: we change the test statistic to avoid overshoot. Our technique uniformly improves power-one SPRTs $(\beta=0)$ for simple nulls and alternatives, or for one-sided nulls and alternatives in exponential families. When $\beta>0$, our techniques provide valid type I error guarantees, lead to similar type II error as Wald’s, but often needs less samples. These improved sequential tests can also be used for deriving tighter parametric confidence sequences, and can be extended to nontrivial settings like sampling without replacement and conformal martingales.
@@ 1 Introduction
Outline of the paper.
In this paper, we propose a general way to uniformly improve (super)martingale based sequential tests by avoiding an overshoot at $1/\alpha$ (Section 2). This means that our sequential tests are never worse than the initial tests, but may stop earlier or are more likely to achieve a rejection. In Section 3, we use this to uniformly improve power-one SPRTs for simple and one-sided nulls and alternatives. We provide nontrivial extensions and applications of our method in Section 4. In Section 5, we demonstrate how our approach should be applied if a stop for futility is to be included. The R code for reproducing the simulation results and calculating the boosting factors is provided in the GitHub repository https://github.com/fischer23/boosting_SPRT.
@@ 2 Improving test martingales by avoiding overshoot
The individual factors $L_{t}$ can be interpreted as sequential e-values recently explored in the literature [18, 31, 7]. A similar (but non-sequential) truncation approach like (8) was proposed by Wang and Ramdas [38] in the context of multiple testing.
Remark 2.
In general we can choose any “boosting function” $\psi_{t}:[0,\infty]\to[0,\infty]$ such that
@@
In order to obtain uniform improvements compared to using the initial factors $L_{t}$, it should be restricted to $\psi_{t}$ with $\psi_{t}(x)\geq x$.
@@ 3 Improving power-one SPRTs
The boosting approach saves on average $4\%$ to $12\%$ of the sample size compared to the SPRT in the considered scenarios, while the percentage increases with the strength of the signal. Furthermore, the boosted SPRT exhausts the significance level, while the SPRT is becoming increasingly conservative for stronger signals.
Figure 1: Comparison of sample size and type I error between the SPRT and the boosted SPRT in a simple Gaussian testing setup for different signal strengths. The boosted SPRT saves $4-12\%$ of sample size compared to the SPRT and exhausts the significance level. The SPRT is becoming increasingly conservative for stronger signals.
@@ 3.2 Handling composite alternatives with the predictable plugin method
There are two approaches that are usually used to handle such composite alternatives in sequential testing. The method of mixtures [19] and the predictable plugin [20, 21]. In the following, we will recap the latter, since it is better suited for our boosting technique.
@@
Robbins and Siegmund [21] introduced different estimators for the case of one-parameter exponential families and evaluated the expected sample size of the resulting sequential tests $\tau_{\text{plugin}}=\inf\{t\in\mathbb{N}:\Lambda_{t}^{\text{plugin}}\geq 1/\alpha\}$. A simple estimator they considered is $\theta_{1}=\theta_{0}$ and
$\theta_{t}=\max\left(\bar{X}_{t-1},\theta_{0}\right),$ (17)
where $\bar{X}_{t-1}=\sum_{i=1}^{t-1}X_{i}$. In Figure 2 we repeated the experiment from Figure 1 in Section 3.1 but assumed that the alternative is not specified in advance using the predictable parameter in (17). Unsurprisingly, the sample size required by both methods is larger than when the alternative was prespecified. However, the gain in sample size achieved by our boosting method is even larger in the case of the composite alternative, as we were able to save $6\%$ to $14\%$ of the sample size compared to the usual SPRT.
@@ 4.1 Confidence sequences
In order to ensure that the resulting CS has this form, it is required that the martingales $(M_{t}^{\theta})_{\theta\in\Theta}$ are nonincreasing in $\theta$ for every $t\in\mathbb{N}$. Now note that each boosting factor $b_{t}^{\theta}$, $\theta\in\Theta$ and $t\in\mathbb{N}$, defined by Algorithm 1, is a nondecreasing function of the (boosted) supermartingale at the previous step $M_{t-1}^{\theta,\mathrm{boost}}$. Therefore, the boosting factors $(b_{t}^{\theta})_{\theta\in\Theta}$ are also nonincreasing in $\theta$. This implies that also the boosted processes $(M_{t}^{\theta,\mathrm{boost}})_{\theta\in\Theta}$ are nonincreasing in $\theta$ for each $t$ such that the resulting CS provides indeed a lower bound. With this, the boosted CS can be easily determined computationally by setting $l_{t}$ at each step $t\in\mathbb{N}$ to be the smallest $\theta\in\Theta$ such that $M_{t}^{\theta,\mathrm{boost}}<1/\alpha$. In the following, we illustrate the construction of a boosted CS with a Gaussian example.
@@
$H_{0}:\mu^{*}=\mu.$
Let $\mu\in\mathbb{R}$ be fixed and define $Z_{t}=X_{t}-\mu$ for all $t\in\mathbb{N}$ such that each $Z_{t}$ is standard normally distributed under $H_{0}$. This implies that
@@
provides a test martingale for $H_{0}$. Furthermore, note that $M_{t}^{\mu}$ is decreasing in $\mu$ for each $t$, such that the family of test martingales $((M_{t}^{\mu}))_{\mu\in\mathbb{R}}$ can be used to find a lower bound $l_{t}$ for $\mu^{*}$. Indeed, we can calculate
@@
This CS was previously proposed by Howard et al. [10]. As described above, one can obtain a boosted confidence sequence numerically by boosting each $(M_{t}^{\mu})$, $\mu\in\mathbb{R}$, separately and choosing the smallest/largest $\mu\in\mathbb{R}$ such that $M_{t}^{\mu,\mathrm{boost}}<1/\alpha$ for the lower/upper bound. In Figure 3 we compare the $(1-\alpha)$ lower confidence bound proposed by Howard et al. [10] with our boosted improvement for $\mu^{*}=2$ (the bounds were obtained by averaging over $100$ trials), which shows that boosting can tighten the CS (20) significantly. We chose $\delta=\sqrt{8\log(1/\alpha)/n}$ for $n=50$, as recommended by Howard et al. [10] for obtaining a tight CS at time $n$.
Note that Howard et al. [10] proved that the CS (20) is valid for all 1-sub-Gaussian distributions. However, our boosting approach only ensures valid error control for Gaussian distributions with variance one. The reason is that
@@
for $b_{t}>1$ and therefore it is not possible to handle all 1-sub-Gaussian distributions simultaneously. It is interesting that it is possible to exploit the information of the observations being exactly Gaussian (instead of just 1-sub-Gaussian), since previous works on confidence sequences did not make use of it [10, 11].
@@ 4.3 Conformal martingales
5 Avoiding conservatism with two-sided thresholds
While the previous sections focused on power-one SPRTs where we only stop to reject the null but not to accept it, we now consider the case where we also wish to stop if it is unlikely that we would reject the null if we continued testing. We will show how we can use a similar boosting technique as in Section 2 to construct powerful anytime-valid tests that allow to stop for futility.
Suppose our test is based on the test (super)martingale $(M_{t})$. We consider stopping for futility (and accepting the null) if $M_{t}\leq\nu_{t}$, where $\nu_{t}<1/\alpha$ is a predictable parameter, meaning $\nu_{t}$ is measurable with respect to $\mathcal{F}_{t-1}$. For example, Wald’s approximation suggests to set $\nu_{t}=\frac{\beta}{1-\alpha}$. Note that stopping for futility can never increase the type I error, but it may hurt the power. Below, we will show how we can gain back part of that power.
@@ 5 Avoiding conservatism with two-sided thresholds
Note that in contrast to Theorem 2.1, Corollary 5.1 does not show that the boosted test $(\delta_{\mathrm{boost}},\tau_{\mathrm{boost}})$ stops earlier than the initial test $(\delta_{M},\tau_{M})$. The problem is that the boosted process is always larger than the initial test until a decision is obtained and therefore may continue sampling while the initial test has already stopped for futility. But, the boosted process never stops later for rejecting the null hypothesis. Therefore, one could obtain a uniform improvement $(\delta^{\prime},\tau^{\prime})$ of $(\delta_{M},\tau_{M})$, meaning $\delta^{\prime}\geq\delta$ and $\tau^{\prime}\leq\tau$ almost surely, by $\delta^{\prime}=\mathbbm{1}\{M_{\tau^{\prime}}^{\mathrm{boost}}\geq 1/\alpha\}$ where $\tau^{\prime}=\min(\tau_{M},\tau_{\mathrm{boost}})$. However, the test $(\delta^{\prime},\tau^{\prime})$ might stop while the boosted supermartingale is strictly positive, meaning it is conservative.
@@
There are many choices for $\nu_{t}$ one could make. For example, one could choose $\nu_{t}=\frac{\beta}{1-\alpha}\prod_{i=1}^{t-1}b_{i}$, where the $b_{i}$ are the previous boosting factors and $\beta\in(0,1)$ is the desired type II error probability. In this case, the true type II error probability is similar to the one obtained by the SPRT with Wald’s approximations, since $\Lambda_{t}^{\mathrm{boost}}\leq\nu_{t}\implies\Lambda_{t}\leq\frac{\beta}{1-\alpha}$. In Figure 5, we compare the sample size, the type II and the type I error probability of the SPRT with our boosted method for different parameters $\beta$ in the same Gaussian simulation setup as considered in Section 3.1 (the type I error probability was obtained via importance sampling, see (15)). The desired type I error probability was set to $\alpha=0.05$ and the strength of the signal to $\delta=0.3$. Note that the boosted method leads to a lower type II error in all cases. Furthermore, it reduces the sample size when $\beta$ is small ($<0.2$), which is the usual case in most applications. For $\beta\geq 0.2$, the sample size was smaller using the SPRT. However, in these cases the boosted method led to a significantly smaller type II error probability. Hence, the sample size obtained with the boosted method could possibly be reduced by choosing a larger stop for futility $\nu_{t}$, while retaining a similar type II error probability as the SPRT. The boosted SPRT approximately exhausted the type I error probability while the SPRT is conservative.
Figure 5: Comparison of sample size, type II and type I error probability between the SPRT and the boosted SPRT in a simple Gaussian testing setup for different desired type II error probabilities ($\beta$). The boosted method saves sample size when $\beta$ is small while reducing the type II error probability. Furthermore, the boosted SPRT approximately exhausts the type I error level $\alpha=0.05$.
@@
If we choose $\gamma_{0}=\beta$ and $\gamma_{1}=1/\alpha$, then the SPRT provably controls the type I and type II error [34, 35] (but is usually conservative). Hence, if we apply our boosted SPRT with $\nu_{t}=\beta\prod_{i=1}^{t-1}b_{i}$, we also control the type I and type II error, since we control the type I error by construction and are uniformly more powerful than the SPRT with $\gamma_{0}=\beta$ and $\gamma_{1}=1/\alpha$.

### v2

@@ Abstract
The sequential probability ratio test (SPRT) by Wald (1945) is a cornerstone of sequential analysis. Based on desired type-I, II error levels $\alpha,\beta\in(0,1)$, it stops when the likelihood ratio statistic crosses certain upper and lower thresholds, guaranteeing optimality of the expected sample size. However, these thresholds are not closed form and the test is often applied with approximate thresholds $(1-\beta)/\alpha$ and $\beta/(1-\alpha)$ (approximate SPRT). When $\beta>0$, this neither guarantees type I,II error control at $\alpha,\beta$ nor optimality. When $\beta=0$ (power-one SPRT), it guarantees type I error control at $\alpha$ that is in general conservative, and thus not optimal. The looseness in both cases is caused by overshoot: the test statistic overshoots the thresholds at the stopping time. One standard way to address this is to calculate the right thresholds numerically, but many papers and software packages do not do this. In this paper, we describe a different way to improve the approximate SPRT: we change the test statistic to avoid overshoot. Our technique uniformly improves power-one SPRTs $(\beta=0)$ for simple nulls and alternatives, or for one-sided nulls and alternatives in exponential families. When $\beta>0$, our techniques provide valid type I and type II error guarantees, while needing less samples than Wald’s approximated thresholds in all considered simulations. These improved sequential tests can also be used for deriving tighter parametric confidence sequences, and can be extended to nontrivial settings like sampling without replacement and conformal martingales.
@@ 1 Introduction
Outline of the paper and our contribution.
In this paper, we propose a general boosting approach to uniformly improve (super)martingale based sequential tests by avoiding an overshoot at $1/\alpha$ (Section 2). This means that our boosted sequential tests are never worse than the initial tests, but may stop earlier or are more likely to achieve a rejection. In Section 3, we use this to uniformly improve power-one SPRTs for simple and one-sided nulls and alternatives. We provide nontrivial extensions and applications of our method in Section 4. In Section 5, we demonstrate how our boosting approach should be applied if a stop for futility is to be included and type II error control is desired. The R code for reproducing the simulation results and calculating the boosting factors is provided in the GitHub repository https://github.com/fischer23/boosting_SPRT.
In Table 1 we compare Wald’s approximate power-one and two-sided SPRT with our boosted versions. In the power-one case, our SPRT provides the same guarantees as Wald’s approximated SPRT, but uniformly improves it in terms of sample size. For the two-sided case, our boosted SPRT provides valid type I and type II error control while Wald’s approximated SPRT does not. In addition, our two-sided SPRT outperformed Wald’s with respect to sample size in all considered simulation scenarios.
Sequential test Type I control Type II control Sample size
Wald’s approx. power-one SPRT power-one Not optimal
Our boosted power-one SPRT power-one Not optimal, but never stops later than Wald’s approx. SPRT
Wald’s approx. two-sided SPRT - - Not optimal
Our boosted two-sided SPRT Not optimal, but often stops earlier than Wald’s approx. SPRT
Table 1: Comparison of the guarantees provided by Wald’s approximate power-one and two-sided SPRT, to our boosted SPRTs. Our boosted SPRT uniformly improves Wald’s approximated SPRT in terms of sample size in the power-one case, and often needs less samples in the two-sided case. In addition, our two-sided SPRT yields valid type I and type II error guarantee, while Wald’s approximated SPRT yields neither.
@@ 2 Improving test martingales by avoiding overshoot
The individual factors $L_{t}$ can be interpreted as sequential e-values recently explored in the literature [20, 33, 8]. A similar (but non-sequential) truncation approach like (8) was proposed by Wang and Ramdas [40] in the context of multiple testing. They also mentioned that in general we can use any “boosting function” $\psi_{t}:[0,\infty]\to[0,\infty]$ such that
@@
However, recently Koning [16] proved that $\psi_{t}(x)=b_{t}x$, where $b_{t}$ is chosen such that (10) is satisfied with an equality, is the log-optimal choice. Therefore, we will stick to the boosting factor for the rest of the paper.
@@ 3 Improving power-one SPRTs
The boosting approach saves on average $4\%$ to $12\%$ of the sample size compared to the SPRT in the considered scenarios, while the percentage increases with the strength of the signal. Furthermore, the boosted SPRT exhausts the significance level, while the SPRT is becoming increasingly conservative for stronger signals. In Appendix A we repeat this experiment for $\alpha=0.01$ instead of $\alpha=0.05$, which yields a similar result.
Figure 1: Comparison of sample size and type I error between the SPRT and the boosted SPRT in a simple Gaussian testing setup for different signal strengths and $\alpha=0.05$. The boosted SPRT saves $4-12\%$ of sample size compared to the SPRT and exhausts the significance level. The SPRT is becoming increasingly conservative for stronger signals.
@@ 3.2 Handling composite alternatives with the predictable plugin method
There are two approaches that are usually used to handle such composite alternatives in sequential testing. The method of mixtures [21] and the predictable plugin [36, 22, 23]. In the following, we will recap the latter, since it is better suited for our boosting technique.
@@
A simple estimator used for the predictable plugin method is the smoothed maximum likelihood estimator (MLE)
$\theta_{t}=\max\left(\frac{\theta_{0}+\sum_{i=1}^{t-1}X_{i}}{t},\theta_{0}\right).$ (17)
In Figure 2 we repeated the experiment from Figure 1 in Section 3.1 but assumed that the alternative is not specified in advance using the predictable parameter in (17). Unsurprisingly, the sample size required by both methods is larger than when the alternative was prespecified. However, the gain in sample size achieved by our boosting method is even larger in the case of the composite alternative, as we were able to save $6\%$ to $14\%$ of the sample size compared to the usual SPRT.
@@ 4.1 Confidence sequences
In order to ensure that the resulting CS has this form, it is required that the martingales $(M_{t}^{\theta})_{\theta\in\Theta}$ are nonincreasing in $\theta$ for every $t\in\mathbb{N}$. Now note that each boosting factor $b_{t}^{\theta}$, $\theta\in\Theta$ and $t\in\mathbb{N}$, defined by Algorithm 1, is a nondecreasing function of the (boosted) supermartingale at the previous step $M_{t-1}^{\theta,\mathrm{boost}}$. Therefore, the boosting factors $(b_{t}^{\theta})_{\theta\in\Theta}$ are also nonincreasing in $\theta$, provided that the distribution of $(M_{t}^{\theta})$ under $H_{0}^{\theta}$ is the same for all $\theta\in\Theta$. This implies that also the boosted processes $(M_{t}^{\theta,\mathrm{boost}})_{\theta\in\Theta}$ are nonincreasing in $\theta$ for each $t$ such that the resulting CS provides indeed a lower bound. With this, the boosted CS can be easily determined computationally by setting $l_{t}$ at each step $t\in\mathbb{N}$ to be the smallest $\theta\in\Theta$ such that $M_{t}^{\theta,\mathrm{boost}}<1/\alpha$. In the following, we illustrate the construction of a boosted CS with a Gaussian example.
@@
$H_{0}^{\mu}:\mu^{*}=\mu.$
Let $\mu\in\mathbb{R}$ be fixed and define $Z_{t}=X_{t}-\mu$ for all $t\in\mathbb{N}$ such that each $Z_{t}$ is standard normally distributed under $H_{0}^{\mu}$. This implies that
@@
provides a test martingale for $H_{0}^{\mu}$. Furthermore, note that $M_{t}^{\mu}$ is decreasing in $\mu$ for each $t$, such that the family of test martingales $((M_{t}^{\mu}))_{\mu\in\mathbb{R}}$ can be used to find a lower bound $l_{t}$ for $\mu^{*}$. Indeed, we can calculate
@@
This CS was already proposed by Robbins and colleagues [6, 21]. As described above, one can obtain a boosted confidence sequence numerically by boosting each $(M_{t}^{\mu})$, $\mu\in\mathbb{R}$, separately and choosing the smallest/largest $\mu\in\mathbb{R}$ such that $M_{t}^{\mu,\mathrm{boost}}<1/\alpha$ for the lower/upper bound. In Figure 3 we compare the $(1-\alpha)$ lower confidence bound proposed by Robbins with our boosted improvement for $\mu^{*}=2$ (the bounds were obtained by averaging over $100$ trials), which shows that boosting can tighten the CS (20) significantly. We chose $\delta=\sqrt{8\log(1/\alpha)/n}$ for $n=50$, as recommended for obtaining a tight CS at time $n$ [11].
Note that Robbins proved that the CS (20) is valid for all 1-sub-Gaussian distributions. However, our boosting approach only ensures valid error control for Gaussian distributions with variance one. The reason is that
@@
for $b_{t}>1$ and therefore it is not possible to handle all 1-sub-Gaussian distributions simultaneously. It is interesting that it is possible to exploit the information of the observations being exactly Gaussian (instead of just 1-sub-Gaussian), since previous works on confidence sequences did not make use of it [6, 21, 11, 12].
@@ 4.3 Conformal martingales
5 Challenging Wald’s two-sided approximate SPRT
While the previous sections focused on power-one SPRTs ($\beta=0$) where we only stop to reject the null but not to accept it, we now consider the case of $\beta>0$, where we also wish to stop if it is unlikely that we would reject the null if we continued testing. We will show how we can use a similar boosting technique as in Section 2 to construct powerful anytime-valid tests that allow to “stop for futility” (stop and accept the null).
Our main comparison point will be Wald’s approximate two-sided SPRT, which uses thresholds of $(1-\beta)/\alpha$ and $\beta/(1-\alpha)$. Importantly, out of the three desiderata of (a) type-I error control at $\alpha$, (b) type-II error control at $\beta$, (c) optimal sample size, these approximate thresholds guarantee none of the three properties. In this section, we will first develop a method with valid type-I error control (as a warm-up), and then one with valid type-I and type-II error control. The latter method often outperforms Wald’s approximate SPRT in the sense that while guaranteeing (a) and (b), our method also manages to typically/often stop sooner than Wald’s (though not guaranteeing (c)).
5.1 Two-sided stopping with valid Type-I error control
Suppose our test is based on a test (super)martingale $(M_{t})$, such as the likelihood ratio. We consider stopping for futility if $M_{t}\leq\nu_{t}$, where $\nu_{t}<1/\alpha$ is a predictable parameter, meaning $\nu_{t}$ is measurable with respect to $\mathcal{F}_{t-1}$. For example, Wald’s approximation suggests to set $\nu_{t}=\frac{\beta}{1-\alpha}$. Note that stopping for futility can never increase the type I error, but it may hurt the power. Below, we will show how we can gain back part of that power.
@@ 5.1 Two-sided stopping with valid Type-I error control
Typically, the parameter $\nu_{t}$ is set to some value smaller than $1$, as $\Lambda_{t}<1$ implies that the current data favors the null hypothesis over the alternative. However, $\nu_{t}$ is a predictable parameter and one could also think of choosing it based on the current value of $\Lambda_{t-1}^{\mathrm{boost}}$. For example, it might be reasonable to set $\nu_{t}=8$ if $\Lambda_{t-1}^{\mathrm{boost}}=10$ to stop in case of decreasing evidence. In general, it is even possible to set $\nu_{t}>\Lambda_{t-1}^{\mathrm{boost}}$.
5.2 Two-sided stopping with valid type-I and type-II error control
We know that boosting always ensures type I error control at the level $\alpha$ by Ville’s inequality (5). Now we want to use the same approach to choose the stop for futility $\nu_{t}$ in a way that ensures valid type II error control at a level $\beta\in(0,1)$. The idea is to use a second SPRT for testing the alternative $H_{1}:X_{i}\sim\mathbb{P}_{1}$ against the null hypothesis $H_{0}:X_{i}\sim\mathbb{P}_{0}$ and stop for futility with our original test if this “inverse SPRT” rejects $H_{1}$. Hence, instead of using the likelihood ratio process $(\Lambda_{t})$ of $\mathbb{P}_{1}$ against $\mathbb{P}_{0}$, we use its inverse $\Lambda_{t}^{\mathrm{inv}}=1/\Lambda_{t}$. By Ville’s inequality (5), we control the type II error for the original test of $H_{0}$ against $H_{1}$ at level $\beta$, if we stop for futility when the inverse likelihood ratio $\Lambda_{t}^{\mathrm{inv}}$ crosses $1/\beta$. However, as we demonstrated in the previous sections, Ville’s inequality is loose if there is overshoot. Hence, to get tighter type II error control, we need to boost the inverse SPRT in the same manner as we did for the usual SPRT and include the stop for futility yielded by the inverse SPRT in the truncation function (22). This requires solving the following two-dimensional equation system at each step $t\in\mathbb{N}$ for the boosting factors $b_{t}$ and $b_{t}^{\mathrm{inv}}$,
[…]
with $\lambda_{t}(X_{t})=p_{1}(X_{t})/p_{0}(X_{t})$,
$\Lambda_{t}^{\mathrm{boost}}=\prod_{i=1}^{t}T_{\alpha}(b_{i}\lambda_{i}(X_{i});\Lambda_{i-1}^{\mathrm{boost}},\nu_{i})\quad\text{and}\quad\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}=\prod_{i=1}^{t}T_{\alpha}(b_{i}/\lambda_{i}(X_{i});\Lambda_{i-1}^{\mathrm{boost},\mathrm{inv}},\nu_{i}^{\mathrm{inv}}).$
The definition of $\nu_{t}$ ensures that the original SPRT only stops for futility when the boosted inverse SPRT stops for rejection. In the following proposition, we use this to prove that this indeed yields valid error control.
Proposition 5.2.
Let the boosting factors $b_{1},b_{2},\ldots$ and inverse boosting factors $b_{1}^{\mathrm{inv}},b_{2}^{\mathrm{inv}},\ldots$ be chosen such that (23) and (24) are fulfilled for each $t\in\mathbb{N}$. Then the boosted SPRT given by $\delta^{\mathrm{boost}}=\mathbbm{1}\{\Lambda_{\tau_{\mathrm{boost}}}^{\mathrm{boost}}\geq 1/\alpha\}$ with $\tau_{\mathrm{boost}}=\inf\{t\in\mathbb{N}:\Lambda_{t}^{\mathrm{boost}}\geq 1/\alpha\lor\Lambda_{t}^{\mathrm{boost}}\leq\nu_{t}\}$ controls the type I error at level $\alpha$ and the type II error at level $\beta$.
Proof.
Due to (23), the type I error control follows immediately by Ville’s inequality. Now, let $\tau_{\mathrm{inv}}=\inf\{t\in\mathbb{N}:\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta\lor\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\leq\nu_{t}^{\mathrm{inv}}\}$ be the stopping time of the inverse SPRT. To prove type II error control, we first show that $\tau_{\mathrm{boost}}=\tau_{\mathrm{inv}}$ almost surely. For this, note that
$\beta\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\alpha\Leftrightarrow\alpha\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\beta,$
and therefore $\nu_{t}<1/\alpha$ iff $\nu_{t}^{\mathrm{inv}}<1/\beta$. On the event $\nu_{1}<1/\alpha,\ldots,\nu_{t}<1/\alpha$, we have
$\Lambda_{t}^{\mathrm{boost}}\leq\nu_{t}\ \Leftrightarrow\ \Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta\quad\text{and}\quad\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\leq\nu_{t}^{\mathrm{inv}}\ \Leftrightarrow\ \Lambda_{t}^{\mathrm{boost}}\geq 1/\alpha.$
Furthermore, if $\nu_{t}=1/\alpha$ and we did not stop until step $t$ ($\tau_{\mathrm{boost}}\geq t$), we always have $\tau_{\mathrm{boost}}=t$ and $\tau_{\mathrm{inv}}=t$. Therefore, $\tau_{\mathrm{boost}}=\tau_{\mathrm{inv}}$. With this, it follows that
$\Lambda_{t}^{\mathrm{boost}}\leq\nu_{t}\land\tau_{\mathrm{boost}}=t$ $\implies\Lambda_{t}\leq\beta\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\land\tau_{\mathrm{boost}}=t$
$\Leftrightarrow\Lambda_{t}^{\mathrm{inv}}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\geq 1/\beta\land\tau_{\mathrm{boost}}=t$
$\Leftrightarrow\Lambda_{t}^{\mathrm{inv}}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\geq 1/\beta\land\tau_{\mathrm{inv}}=t$
$\Leftrightarrow\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta\land\tau_{\mathrm{inv}}=t,$
where the last equivalence follows from the fact that if we have not stopped for futility yet, then $\Lambda_{t}^{\mathrm{inv}}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}\geq 1/\beta$ iff $\Lambda_{t}^{\mathrm{boost},\mathrm{inv}}\geq 1/\beta$. Due to Ville’s inequality and (24), the probability that there exists a $t\in\mathbb{N}$ satisfying the latter inequality is bounded by $\beta$ under $H_{1}$. ∎
Note that the error control of this strategy can still be slightly conservative, since it can happen that
$\beta\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\alpha\Leftrightarrow\alpha\prod_{i=1}^{t}b_{i}\prod_{i=1}^{t}b_{i}^{\mathrm{inv}}>1/\beta,$
which means that the rejection region of the boosted SPRT and the boosted inverse SPRT overlap. For this reason, we set $\nu_{t}=1/\alpha$ and $\nu_{t}^{\mathrm{inv}}=1/\beta$ in this case which ensures valid type I and type II error control but avoids the simultaneous acceptance and rejection of both hypotheses. However, this conservatism is expected to be small, particularly if $\alpha$ and $\beta$ are small.
Figure 5: Comparison of sample size, type II and type I error probability between the SPRT with Wald’s approximated thresholds and the boosted SPRT in a simple Gaussian testing setup for different desired type II error probabilities ($\beta$) and $\alpha=0.01$. The boosted method provably controls the type I and type II error, while the SPRT with Wald’s approximated thresholds does not. Furthermore, the boosted SPRT saves sample size and nearly exhausts the error probabilities.
In Figure 5, we compare the sample size, the type II and the type I error probability of the SPRT with Wald’s approximated thresholds to our boosted method for different parameters $\beta$ in the same Gaussian simulation setup as considered in Section 3.1 (the type I error probability was obtained via importance sampling, see (15)). The desired type I error probability was set to $\alpha=0.05$ and the strength of the signal to $\delta=0.3$. The boosted method reduced the required sample size in all cases. Furthermore, the boosted SPRT nearly exhausts the type I error and type II error probability, particularly if $\beta$ is small, while the SPRT is conservative. Also note that the boosted SPRT provably controls the type I and type II error, but the SPRT with Wald’s approximated thresholds does not. In Appendix A we repeat this experiment for $\alpha=0.01$ instead of $\alpha=0.05$, which leads to similar results.
Remark 5.
In order to avoid solving the system of equations (23) and (24), one could also just set $\nu_{t}=\min\left(\beta\prod_{i=1}^{t}b_{i},1/\alpha\right)$ and solve for $b_{t}$ in
$\mathbb{E}_{0}\left[T_{\alpha}(b_{t}\lambda_{t}(X_{t});\Lambda_{t-1}^{\mathrm{boost}},\nu_{t})\Big{|}\mathcal{F}_{t-1}\right]=1.$
This would still lead to valid type I and type II error control, however, the type II error control would usually be more conservative than with the proposed approach.
Remark 6.
To ensure error control it is generally sufficient to satisfy (23) and (24) with inequalities $\leq$ (for all four equations). However, in order to obtain tight control the parameters $b_{t}$ and $b_{t}^{\mathrm{inv}}$ should be chosen as large as possible.